OpenStudy (anonymous):

Find a ⋅ b. a = <2, 4>, b = <2, 5>

OpenStudy (anonymous):

<4, 9> 24 <4, 20> 16

jimthompson5910 (jim_thompson5910):

Similar Problem Given a = <7,19> and b = <21,8> find a ⋅ b In general, if a = <x,y> and b = <z,w> then a ⋅ b = x*z + y*w so a ⋅ b = 7*21 + 19*8 a ⋅ b = 147 + 152 a ⋅ b = 299

OpenStudy (anonymous):

thank you that help alot